OpenStudy (anonymous):

how do I write equation for circle with diameter having endpoints (-3,4) and (5,-2)

OpenStudy (anonymous):

Find midpoint of the diameter with these two given points first..

OpenStudy (anonymous):

and once I find the midpoint how do I find radius?

OpenStudy (anonymous):

\[Midpoint (x,y) = (\frac{x_1 + x_2}{2}, \frac{y_1 + y_2}{2})\]

OpenStudy (anonymous):

is the midpoint (1,1)

OpenStudy (xapproachesinfinity):

distance formula for radius

OpenStudy (anonymous):

For Radius, first find the distance..

OpenStudy (anonymous):

radius 5??

OpenStudy (anonymous):

That will given you the diameter length, for radius, divide the distance by 2..

OpenStudy (anonymous):

Yeah you are pretty much right.. :)

OpenStudy (anonymous):

what do you mean pretty much. that doesn't count in math

OpenStudy (anonymous):

Centrre (1,1) and Radius = 5.. Now use Standard Equation of Circle to write equation..

OpenStudy (anonymous):

(x-1)^2 +(y-1)^2=25

OpenStudy (anonymous):

Great... :)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Welcome.. :)